# FINALLY: fast and universal speech enhancement with studio-like quality

### v1

@@ Abstract
In this paper, we address the challenge of speech enhancement in real-world recordings, which often contain various forms of distortion, such as background noise, reverberation, and microphone artefacts. We revisit the use of Generative Adversarial Networks (GANs) for speech enhancement and theoretically show that GANs are naturally inclined to seek the point of maximum density within the conditional clean speech distribution, which, as we argue, is essential for the speech enhancement task. We study various feature extractors for perceptual loss to facilitate the stability of adversarial training, developing a methodology for probing the structure of the feature space. This leads us to integrate WavLM-based perceptual loss into MS-STFT adversarial training pipeline, creating an effective and stable training procedure for the speech enhancement model. The resulting speech enhancement model, which we refer to as FINALLY, builds upon the HiFi++ architecture, augmented with a WavLM encoder and a novel training pipeline. Empirical results on various datasets confirm our model’s ability to produce clear, high-quality speech at 48 kHz, achieving state-of-the-art performance in the field of speech enhancement.
@@ 3 Perceptual Loss for Speech Generation
One promising idea to reduce the averaging effect is to choose an appropriate representation space for regression, which is less “entangled” than waveform or spectrogram space. In simpler terms, the regression space should be designed so that averaged representation of sounds that are indistinguishable to humans (such as the same phoneme spoken by the same speaker with the same prosody) is still representation of this sound (see Section A.1).
@@
As can be seen, the features extracted by the convolutional encoder of the WavLM model present the best performance according to both our internal metrics and MOS score on the neural vocoding task. We have also tested the other layers of WavLM but did not observe significant improvements when using other layers (see Section A.4). Empirically, we have found that if WavLM-conv feature loss is used together with spectrogram loss (L1-distance magnitudes of STFT), the performance on the vocoding task increases significantly, likely due to the looseness of the WavLM representation (Section A.1). Without any adversarial training, the WavLM-conv+L1-STFT loss (LMOS-loss, Equation 2) achieves an MOS score of $4.31\pm 0.08$, approaching the original adversarially trained HiFi GAN generator which achieves $4.69\pm 0.05$ (Section A.3).
@@ 5 Related Work
Koizumi et al. (2023c) proposes to use w2v-BERT features (Chung et al., 2021) as intermediate representations for speech restoration. The features extracted from the noisy waveform are processed by a feature cleanser which is conditioned on the text representation extracted from transcripts via PnG-BERT, and speaker embedding extracted from the waveform. The feature cleanser is trained to minimize a regression loss between w2v-BERT features extracted from the clean signal and the predicted ones. At the second stage, the WaveFit vocoder is trained to synthesize a waveform based on the predicted features (Koizumi et al., 2023a). An important difference with our work is that the method uses a text transcript, while our model does not. Despite this difference, we show that our model delivers better perceptual quality as reported by human listeners (Appendix B).
@@ 6.1 Evaluation
UNIVERSE Data: 100 audio clips randomly generated by the UNIVERSE (Serrà et al., 2022) authors from clean utterances sampled from VCTK and Harvard sentences, together with noises/backgrounds from DEMAND and FSDnoisy18k. The data contains various artificially simulated distortions including band limiting, reverberation, codec, and transmission artefacts. Please refer to (Serrà et al., 2022) for further details.
The validation data of UNIVERSE was artificially simulated from clean speech recordings using the same pipeline that the authors utilized for training. Therefore, we must note that the comparison is conducted in a manner advantageous to UNIVERSE, since our data simulation pipeline is different.
@@
In addition, we compute the Phoneme Error Rate (PhER) and Word Error Rate (WER) by comparing ground truths with the generated samples (see Appendix D for details) for both the UNIVERSE and VCTK-DEMAND datasets. For subjective quality assessment, we conduct 5-point Mean Opinion Score (MOS) tests. All audio clips are normalized to ensure volume differences do not influence the raters’ evaluations. The raters are required to be English speakers using appropriate listening equipment (more details in Figure 10).
@@
Further, we evaluate our model on the VCTK-DEMAND dataset (Valentini-Botinhao et al., 2017) with additional metrics such as PESQ (Rix et al., 2001a), STOI (Taal et al., 2011), and SI-SDR (Roux et al., 2018). These metrics are included to ensure consistent comparison with previous works.
@@ 6.2 Comparison with existing approaches
We consider BBED (Lay et al., 2023), STORM (Lemercier et al., 2023), and UNIVERSE (Serrà et al., 2022) diffusion models, along with Voicefixer and DEMUCS regression models, as our baselines. In addition, we consider our closest competitor, HiFi-GAN-2, as a GAN-based baseline. The data for comparison with HiFi-GAN-2 and UNIVERSE were taken from their demo pages, since the authors did not release any code. We conduct comparisons with BBED, STORM, Voicefixer, DEMUCS, and HiFi-GAN-2 on real-world VoxCeleb1 samples, and the comparison with UNIVERSE on the simulated data provided by the authors of this work. The results are presented in Table 2. We also compare these models in VCTK-DEMAND dataset, provided in Table 3. We complement this table by two additional models: MetricGAN+ (Fu et al., 2021) and DB-AIAT (Yu et al., 2021).
@@
Lastly, we would like to comment on the results in Table 3. Our model outperforms baselines in subjective evaluation and no-reference metrics, e.g. UTMOS, but significantly in terms of PESQ (Rix et al., 2001a), STOI (Taal et al., 2011) and SI-SDR (Roux et al., 2018). However, there have been numerous works consistently reporting low correlation of reference-based metrics with human perceptual judgment (Manocha et al., 2022; Manjunath, 2009; Andreev et al., 2022). In particular, the study (Manocha et al., 2022) reports that no-reference metrics (including DNSMOS, reported in our work) correlate significantly better with human perception and therefore have higher relevance for objective comparison between methods. Furthermore, in our study, we report the MOS score, which directly reflects human judgments of restoration quality.
@@ 6.3 Ablation study
Firstly, we compare the LMOS loss against two other regression losses in the context of training a small speech enhancement model (10 times smaller than the final model; please see the Section D.2 for details). The first regression loss is the Mel-Spectrogram loss, which was proposed by Kong et al. (2020). As the second alternative, we employ the Reconstruction loss (RecLoss) proposed by Défossez et al. (2023). It consists of a combination of L1 and L2 distances between mel-spectrograms computed with different resolutions. For these experiments, we conducted a grid search for the weights of each reconstruction loss (details are in Section D.2) and report results for the best option in Table 4.
@@
The other proposed improvements bring incremental gains in perceptual MOS quality. Firstly, we add the features of the WavLM encoder as an additional inputs to HiFi++. Next, we scale the architecture by increasing the number of channels within the model. After that, we concatenate the Upsample WaveUNet with the FINALLY-16 architecture and implement the 3rd stage of training on the DAPS-clean dataset (48 kHz). Finally, we use additional human feedback losses (HF Loss) during the 3rd stage to further improve perceptual quality.
Since Table 4 does not clearly demonstrate the advantages of using the WavLM (Chen et al., 2022b) encoder in terms of MOS, we have conducted additional ablation study on the more challenging UNIVERSE (Serrà et al., 2022) validation dataset. The results of these additional experiments, which are presented in Table 5, clearly demontarte the importance of WavLM encoder.
@@ 7 Conclusion
Appendix A Additional results for perceptual losses
A.1 More on motivation
@@
A.2 Influence of losses on final audio
@@
A.3 Comparison with prior work
@@ A.4 WavLM layers
A.5 WavLM features visualization
@@
Appendix B Comparison with MIIPHER
We compare the enhancement quality of LibriTTS test_other samples as released by the Miipher authors (Koizumi et al., 2023b). The results are outlined in Table 7. For the information on WER refer to Section D.1. computations.
@@ Appendix B Comparison with MIIPHER
Appendix C Implementation details
C.1 Reproducibility statement
Our model is based on HiFi++ architecture (Andreev et al., 2022), for which code is available in https://github.com/SamsungLabs/hifi_plusplus. We used several open-source datasets for training. More specifically, we used DAPS, LibriTTS-R and Deep Noise Suppression Challenge (DNS) (only noises). We rely on official implementations of MS-STFT discriminators, which can be found in https://github.com/facebookresearch/encodec. For LMOS loss and HiFi++ conditioining, we used WavLM large model from Hugging Face, which can be found in https://huggingface.co/microsoft/wavlm-large.
C.2 Augmentations
@@
C.3 Model architecture
In this section we detail parameters for our speech enhancement models. We use HiFi++ (Andreev et al., 2022) as backbone. This architecture consists of four parts: Spectral UNet, HiFi Upsampler, WaveUNet and SpectalMaskNet. In addition to that we also use WavLM-large feature encoder with transformer features, which takes waveform as an input and outputs last hidden state of the transformer, which is stacked with the SpectralUNet output and fed into the HiFi Upsampler. In the next paragraphs we thoroughly discuss the architecture of every part. Architecture summary is presented in Table 9
@@ C.3 Model architecture
To better understand Spectral UNet, WaveUNet and SpectralMaskNet we first discuss the key building block. Each residual block is composed of convolution, followed by LeakyReLU. We use weight norm (Salimans & Kingma, 2016) instead of BatchNorm, we found it to be more efficient in our preliminary experiments. We use kernel size 3 for 2D UNets and kernel size 5 for 1D UNet. Each residual block has additive skip connection. In our architectures we use a stacked composition of residual blocks. The number of residual blocks stacked together in one layer is called depth.
@@
SpectralUNet processes a mel-spectrogram input with dimensions [B, 80, T] and outputs a hidden state with dimensions [B, 512, T]. The input mel-spectrogram is combined with positional encoding. The architecture is based on a 2D UNet (Ronneberger et al., 2015) with five layers, each having respective channels of $[16,32,64,128,256]$ and a depth of 4. Additional convolutions are applied after these layers to transform 2D data from shape [B, 16, W, T] to [B, 1, W, T], and then to [B, 512, T]. This output is concatenated along the channel dimension with the final transformer hidden state of WavLM (Chen et al., 2022b), which has been interpolated using the nearest-neighbour method to match the output length of SpectralUNet in the time dimension. The concatenated result then passes through a Residual block with kernel size 3 and with a width of 1536 (1024 from the final transformer hidden state and 512 from SpectralUNet output),than passes through 1d Convolution with kernel size 1 and LeakyReLU to obtain size 512 in the channel dimension, and is subsequently fed into the HiFi Upsampler.
@@
We employ MS-STFT discriminators (Défossez et al., 2023) for adversarial training. We use 5 MS-STFT discriminators with n_ffts $[4096,2048,1024,512,256]$, hop_lengths $[1024,512,256,128,64]$, win_lengths $[4096,2048,1024,512,256]$ in 48kHz setting, for 16kHz one all above mentioned parameters are divided by 2. Our implementation closely follows (Défossez et al., 2023).
@@
The main model is trained using 8 Nvidia P40 GPUs with effective batch size 32 and AdamW (Loshchilov & Hutter, 2017) as our main optimizer. For pretraining we use learning rate 0.0002, betas (0.8, 0.99) and learning rate exponential decay of 0.996 for each 200 iterations, the pretraining lasts 100,000 iterations. For the second stage we use learning rate 0.0002 with betas (0.8, 0.99) and learning rate decay 0.995, whereas for discriminators we use learning rate 0.0002 with betas (0.5, 0.999) and learning rate decay 0.995, the discriminators perform 2 optimization iterations for every 1 generator’s optimization iteration, the training lasts 30,000 generator’s iterations. We also use a linear warm-up for 2000 iterations for the generator. We use the same training parameters for the third stage during generator’s 40,000 iterations. The first and the second stage are trained in 16kHz using LibriTTS-R dataset with noises from DNS and augmentations, discussed in Footnote 2. Whereas the third stage is trained in 48kHz using DAPS dataset with noises from DNS as well. Datasets are discussed in Section C.1
@@
To better contextualize our results we provide a detailed comparison of the model sizes and training data of the baseline models.
@@
C.4 Limitations
The primary area for further developing in the proposed model lies in improving perceptual quality, particularly in instances when the speech is severely degraded. Although our model is capable of capturing the content of such signals, even in situations that may be impossible for humans, some artifacts remain. These artifacts could possibly be attributed to the high uncertainty of the initial signal, such as the speaker’s voice, for instance.

### v2

@@ Abstract
In this paper, we address the challenge of speech enhancement in real-world recordings, which often contain various forms of distortion, such as background noise, reverberation, and microphone artefacts. We revisit the use of Generative Adversarial Networks (GANs) for speech enhancement and theoretically show that GANs are naturally inclined to seek the point of maximum density within the conditional clean speech distribution, which, as we argue, is essential for the speech enhancement task. We study various feature extractors for perceptual loss to facilitate the stability of adversarial training, developing a methodology for probing the structure of the feature space. This leads us to integrate WavLM-based perceptual loss into MS-STFT adversarial training pipeline, creating an effective and stable training procedure for the speech enhancement model. The resulting speech enhancement model, which we refer to as FINALLY, builds upon the HiFi++ architecture, augmented with a WavLM encoder and a novel training pipeline. Empirical results on various datasets confirm our model’s ability to produce clear, high-quality speech at 48 kHz, achieving state-of-the-art performance in the field of speech enhancement. Demo page: https://samsunglabs.github.io/FINALLY-page/
@@ 3 Perceptual Loss for Speech Generation
One promising idea to reduce the averaging effect is to choose an appropriate representation space for regression, which is less “entangled” than waveform or spectrogram space. In simpler terms, the regression space should be designed so that averaged representation of sounds that are indistinguishable to humans (such as the same phoneme spoken by the same speaker with the same prosody) is still representation of this sound (see Section B.1).
@@
As can be seen, the features extracted by the convolutional encoder of the WavLM model present the best performance according to both our internal metrics and MOS score on the neural vocoding task. We have also tested the other layers of WavLM but did not observe significant improvements when using other layers (see Section B.4). Empirically, we have found that if WavLM-conv feature loss is used together with spectrogram loss (L1-distance magnitudes of STFT), the performance on the vocoding task increases significantly, likely due to the looseness of the WavLM representation (Section B.1). Without any adversarial training, the WavLM-conv+L1-STFT loss (LMOS-loss, Equation 2) achieves an MOS score of $4.31\pm 0.08$, approaching the original adversarially trained HiFi GAN generator which achieves $4.69\pm 0.05$ (Section B.3).
@@ 5 Related Work
Koizumi et al. (2023c) proposes to use w2v-BERT features (Chung et al., 2021) as intermediate representations for speech restoration. The features extracted from the noisy waveform are processed by a feature cleanser which is conditioned on the text representation extracted from transcripts via PnG-BERT, and speaker embedding extracted from the waveform. The feature cleanser is trained to minimize a regression loss between w2v-BERT features extracted from the clean signal and the predicted ones. At the second stage, the WaveFit vocoder is trained to synthesize a waveform based on the predicted features (Koizumi et al., 2023a). An important difference with our work is that the method uses a text transcript, while our model does not. Despite this difference, we show that our model delivers better perceptual quality as reported by human listeners (Appendix C).
@@ 6.1 Evaluation
UNIVERSE Data: 100 audio clips randomly generated by the UNIVERSE (Serrà et al., 2022) authors from clean utterances sampled from VCTK and Harvard sentences, together with noises/backgrounds from DEMAND and FSDnoisy18k. The data contains various artificially simulated distortions including band limiting, reverberation, codec, and transmission artefacts. Please refer to (Serrà et al., 2022) for further details. The validation data of UNIVERSE is artificially simulated from clean speech recordings using the same pipeline that the authors utilized for training. Therefore, we must note that the comparison is conducted in a manner advantageous to UNIVERSE, since our data simulation pipeline is different.
@@
In addition, we compute the Phoneme Error Rate (PhER) and Word Error Rate (WER) by comparing ground truths with the generated samples (see Appendix E for details) for both the UNIVERSE and VCTK-DEMAND datasets. For subjective quality assessment, we conduct 5-point Mean Opinion Score (MOS) tests. All audio clips are normalized to ensure volume differences do not influence the raters’ evaluations. The raters are required to be English speakers using appropriate listening equipment (more details in Figure 10).
@@
We also evaluate our model on the VCTK-DEMAND dataset (Valentini-Botinhao et al., 2017) with additional metrics such as PESQ (Rix et al., 2001a), STOI (Taal et al., 2011), and SI-SDR (Roux et al., 2018). These metrics are included to ensure consistent comparison with previous works.
@@ 6.2 Comparison with existing approaches
We consider BBED (Lay et al., 2023), STORM (Lemercier et al., 2023), and UNIVERSE (Serrà et al., 2022) diffusion models, along with Voicefixer and DEMUCS regression models, as our baselines. In addition, we consider our closest competitor, HiFi-GAN-2, as a GAN-based baseline. The data for comparison with HiFi-GAN-2 and UNIVERSE were taken from their demo pages, since the authors did not release any code. We conduct comparisons with BBED, STORM, Voicefixer, DEMUCS, and HiFi-GAN-2 on real-world VoxCeleb1 samples and the comparison with UNIVERSE on the simulated data, provided by the authors of this work. The results are presented in Table 2. We also compare these models on VCTK-DEMAND dataset, results can be found in Table 3. We complement this table by two additional models: MetricGAN+ (Fu et al., 2021) and DB-AIAT (Yu et al., 2021).
@@
Lastly, we would like to comment on the results in Table 3. Our model outperforms baselines in subjective evaluation and no-reference metrics, e.g. UTMOS, but underperforms in terms of PESQ (Rix et al., 2001a), STOI (Taal et al., 2011) and SI-SDR (Roux et al., 2018). Notably, there are numerous works consistently reporting low correlation of reference-based metrics with human perceptual judgment (Manocha et al., 2022; Manjunath, 2009; Andreev et al., 2022). In particular, the study (Manocha et al., 2022) reports that no-reference metrics (including DNSMOS, reported in our work) correlate significantly better with human perception and therefore have higher relevance for objective comparison between methods. Furthermore, in our study, we report the MOS score, which directly reflects human judgments of restoration quality.
@@ 6.3 Ablation study
Firstly, we compare the LMOS loss against two other regression losses in the context of training a small speech enhancement model (10 times smaller than the final model; please see the Section E.2 for details). The first regression loss is the Mel-Spectrogram loss, which was proposed by Kong et al. (2020). As the second alternative, we employ the Reconstruction loss (RecLoss) proposed by Défossez et al. (2023). It consists of a combination of L1 and L2 distances between mel-spectrograms computed with different resolutions. For these experiments, we conducted a grid search for the weights of each reconstruction loss (details are in Section E.2) and report results for the best option in Table 4.
@@
The other proposed improvements bring incremental gains in perceptual MOS quality. Firstly, we add the features of the WavLM encoder as an additional input to HiFi++. Next, we scale the architecture by increasing the number of channels within the model. After that, we concatenate the Upsample WaveUNet with the FINALLY-16 architecture and implement the 3rd stage of training on the DAPS-clean dataset (48 kHz). Finally, we use additional human feedback losses (HF Loss) during the 3rd stage to further improve perceptual quality.
Since Table 4 does not clearly demonstrate the advantages of using the WavLM (Chen et al., 2022b) encoder in terms of MOS, we provide additional ablation study on more challenging UNIVERSE (Serrà et al., 2022) validation dataset. The results of these additional experiments, presented in Table 5, clearly demonstrate the importance of WavLM encoder.
@@ 7 Conclusion
Appendix A Proof
See 1
Proof.
[…]
Since $p_{\text{clean}}(y|x)$ is Lipschitz continuous, there exist $L>0$ and $C>0$ such that for any $\xi>C$ if $y-g_{\theta}(x)\in[-1/\xi,1/\xi]^{n}$, then $p_{\text{clean}}(y|x)\in(p_{\text{clean}}(g_{\theta}(x)|x)-L/\xi,p_{\text{clean}}(g_{\theta}(x)|x)+L/\xi)$.
Therefore, the divergence could be lower-bounded by
$\frac{1}{2}\cdot\frac{(1-\frac{2^{n}}{\xi^{n}}\cdot(p_{\text{clean}}(g_{\theta}(x)|x)+L/\xi))^{2}}{1+\frac{2^{n}}{\xi^{n}}\cdot(p_{\text{clean}}(g_{\theta}(x)|x)+L/\xi)}+\frac{1}{2}-\frac{2^{n-1}}{\xi^{n}}(p_{\text{clean}}(g_{\theta}(x)|x)+L/\xi),$
and upper-bounded by
$\frac{1}{2}\cdot\frac{(1-\frac{2^{n}}{\xi^{n}}\cdot(p_{\text{clean}}(g_{\theta}(x)|x)-L/\xi))^{2}}{1+\frac{2^{n}}{\xi^{n}}\cdot(p_{\text{clean}}(g_{\theta}(x)|x)-L/\xi)}+\frac{1}{2}-\frac{2^{n-1}}{\xi^{n}}(p_{\text{clean}}(g_{\theta}(x)|x)-L/\xi).$
After taking Taylor expansion up to the linear term of $1/\xi^{n}$, we get that
$1-\frac{2^{n+1}}{\xi^{n}}p_{\text{clean}}(g_{\theta}(x)|x)+o(1/\xi^{n})\leq\chi^{2}_{\text{Pearson}}(p_{g}^{\xi}(p_{\text{clean}}+p_{g}^{\xi})/2),$
and
$\chi^{2}_{\text{Pearson}}(p_{g}^{\xi}(p_{\text{clean}}+p_{g}^{\xi})/2)\leq 1-\frac{2^{n+1}}{\xi^{n}}p_{\text{clean}}(g_{\theta}(x)|x)+o(1/\xi^{n}).$
Thus,
$\chi^{2}_{\text{Pearson}}(p_{g}^{\xi}||(p_{\text{clean}}+p_{g}^{\xi})/2)=1-\frac{2^{n+1}}{\xi^{n}}p_{\text{clean}}(g_{\theta}(x)|x)+o(1/\xi^{n}).$
[…]
Appendix B Additional results for perceptual losses
B.1 More on motivation
@@
B.2 Influence of losses on final audio
@@
B.3 Comparison with prior work
@@ B.4 WavLM layers
B.5 WavLM features visualization
@@
Appendix C Comparison with MIIPHER
We compare the enhancement quality of LibriTTS test_other samples as released by the Miipher authors (Koizumi et al., 2023b). The results are outlined in Table 7. For the information on WER refer to Section E.1. computations.
@@ Appendix C Comparison with MIIPHER
Appendix D Implementation details
D.1 Reproducibility statement
Our model is based on HiFi++ architecture (Andreev et al., 2022), for which code is available in https://github.com/SamsungLabs/hifi_plusplus. We used several open-source datasets for training. More specifically, we used DAPS, LibriTTS-R and Deep Noise Suppression Challenge (DNS) (only noises). We rely on official implementations of MS-STFT discriminators, which can be found in https://github.com/facebookresearch/encodec. For LMOS loss and HiFi++ conditioning, we used WavLM large model from Hugging Face, which can be found in https://huggingface.co/microsoft/wavlm-large.
D.2 Augmentations
@@
D.3 Model architecture
In this section we detail parameters for our speech enhancement models. We use HiFi++ (Andreev et al., 2022) as backbone. This architecture consists of four parts: Spectral UNet, HiFi Upsampler, WaveUNet and SpectalMaskNet. In addition to that, we use WavLM-large feature encoder with transformer features, which takes a waveform as an input and outputs the last hidden state of the transformer, which is stacked with the SpectralUNet output and fed into the HiFi Upsampler. In the next paragraphs, we thoroughly discuss the architecture of each component. Architecture summary is presented in Table 9
@@ D.3 Model architecture
To better understand Spectral UNet, WaveUNet and SpectralMaskNet we first discuss the key building block. Each residual block is composed of convolution, followed by LeakyReLU. We use weight norm (Salimans & Kingma, 2016) instead of BatchNorm, we found it to be more efficient in our preliminary experiments. We use kernel size 3 for 2D UNets and kernel size 5 for 1D UNet. Each residual block has additive skip connection. In our architecture, we use a stacked composition of residual blocks. The number of residual blocks stacked together in one layer is called depth.
@@
SpectralUNet processes a mel-spectrogram input with dimensions [B, 80, T] and outputs a hidden state with dimensions [B, 512, T]. The input mel-spectrogram is combined with positional encoding. The architecture is based on a 2D UNet (Ronneberger et al., 2015) with five layers, each having respective channels of $[16,32,64,128,256]$ and a depth of 4. Additional convolutions are applied after these layers to transform 2D data from shape [B, 16, W, T] to [B, 1, W, T], and then to [B, 512, T]. This output is concatenated along the channel dimension with the final transformer hidden state of WavLM (Chen et al., 2022b), which has been interpolated using the nearest-neighbour method to match the output length of SpectralUNet in the time dimension. The concatenated result then passes through a Residual block with kernel size 3 and with a width of 1536 (1024 from the final transformer hidden state and 512 from SpectralUNet output), then passes through 1d Convolution with kernel size 1 and LeakyReLU to obtain size 512 in the channel dimension, and is subsequently fed into the HiFi Upsampler.
@@
We employ MS-STFT discriminators (Défossez et al., 2023) for adversarial training. We use 5 MS-STFT discriminators with n_ffts $[4096,2048,1024,512,256]$, hop_lengths $[1024,512,256,128,64]$, win_lengths $[4096,2048,1024,512,256]$ in 48kHz setting, for 16kHz all above-mentioned parameters are divided by 2. Our implementation closely follows (Défossez et al., 2023).
@@
The main model is trained using 8 Nvidia P40 GPUs with effective batch size 32 and AdamW (Loshchilov & Hutter, 2017) as our main optimizer. For pretraining we use learning rate 0.0002, betas (0.8, 0.99) and learning rate exponential decay of 0.996 for each 200 iterations, the pretraining lasts 100,000 iterations. For the second stage we use learning rate 0.0002 with betas (0.8, 0.99) and learning rate decay 0.995, whereas for discriminators we use learning rate 0.0002 with betas (0.5, 0.999) and learning rate decay 0.995, the discriminators perform 2 optimization iterations for every 1 generator’s optimization iteration, the training lasts 30,000 generator’s iterations. We also use a linear warm-up for 2000 iterations for the generator. We use the same training parameters for the third stage during generator’s 40,000 iterations. The first and the second stage are trained in 16kHz using LibriTTS-R dataset with noises from DNS and augmentations, discussed in Footnote 2. Whereas the third stage is trained in 48kHz using DAPS dataset with noises from DNS as well. Datasets are discussed in Section D.1
@@
To better contextualize our results, we provide a detailed comparison of the model sizes and training data of the baseline models.
@@
D.4 Limitations
The primary area for further developing in the proposed model lies in improving perceptual quality, particularly in instances when the speech is severely degraded. Although our model is capable of capturing the content of such signals, even in situations that may be impossible for humans, some artefacts remain. These artefacts could possibly be attributed to the high uncertainty of the initial signal, such as the speaker’s voice, for instance.